OpenStudy (narissa):

determine if the statement is true or false when evaluated 2x-x/3<_5 when x=3

OpenStudy (anonymous):

hard to read is it \[2x-\frac{x}{3}\leq 5\]?

OpenStudy (narissa):

its true right?

OpenStudy (narissa):

yes

OpenStudy (anonymous):

ok so lets to it \[2\times 3-\frac{3}{3}+?\]

OpenStudy (anonymous):

i meant lets compute \[2\times 3-\frac{3}{3}\]

OpenStudy (narissa):

it would be equal no

OpenStudy (anonymous):

yes i get 5 and \[5\leq 5\] is true

OpenStudy (narissa):

6- one whole is 5?

OpenStudy (anonymous):

yes

OpenStudy (narissa):

thanks ur smart